Consider the labor demand curve given by the equation

                                                                                                  100

                                                                                                    L =  -------

          W

For small changes in the wage around any given point on the curve, the slope of the curve (the change in L divided by the change in W) is given by the formula

                                                                                                  100

                                                                                            Slope = −    --------

                              W 2

a. Find the slope of the curve for a small change in the wage around W = $5.

b. Find the own-wage elasticity of labor demand for a small change in the wage around $5.
Answer #1

A. It is given that slope of the curve is given by= -100/W2. Putting values in, we get

slope=-100/52= -100/25= -4

B. Own wage elasticity of demand is given by

the % change in employment due to a 1% change in the wage rate

In formula terms, this can be given by

(Change in L/Change in w)*(w/L)

Putting values

=-(100/w2)(w/L)

L is given as 100/w. So,

=-(100/w2)(w2/100)

=-1